A buffer contains equal concentrations of a weak acid HA and its conjugate base A−. If the value of Ka for HA is 1.0 × 10−9 , what is the pH of the buffer?

1. 7.0

2. 1.0

3. 5.0

4. 13.0

5. 9.0


1
Expert's answer
2021-03-15T09:07:58-0400


1.0×10−9=[H+][NH3][NH+4]=x(0.0500)

0.0350

5.6×10−10=[H+][NH3][NH4+]=x(0.0500)0.0350

Solving for x (H+):

x = [H+] = 3.92 x 10-10

pH = -log(3.92 x 10-10)

pH = 9.41

= 9